Consider Derek's budget information: materials to be used, $64,800; direct labor, $201,500; factory overhead, $398,800; work in process inventory on January 1, $185,600; and work in progress inventory on December 31, $197,000. What is the budgeted cost of goods manufactured for the year?

a.$653,700

b.$850,700

c.$197,000

d.$665,100

Use the information below for Nuthatch Corporation to answer the question that follow.

Nuthatch Corporation began its operations on September 1 of the current year. Budgeted sales for the first three months of business—September, October, and November—are $236,000, $316,000, and $412,000, respectively. The company expects to sell 30% of its merchandise for cash. Of sales on account, 80% are expected to be collected in the month of the sale and 20% in the month following the sale.

The cash collections expected in September from accounts receivable are estimated to be

a.$165,200

b.$236,000

c.$132,160

d.$283,200

Woodpecker Co. has $310,000 in accounts receivable on January 1. Budgeted sales for January are $934,000. Woodpecker Co. expects to sell 20% of its merchandise for cash. Of the remaining 80% of sales on account, 75% are expected to be collected in the month of sale and the remainder the following month. The January cash collections from sales are

a.$845,760

b.$1,057,200

c.$1,367,200

d.$634,320
Answer #1

1) Cost of goods manufactured

Total manufacturing cost = 64800+201500+398800 = 665100

Cost of goods manufactured = 185600+665100-197000 = 653700

So answer is a) $653700

2) Cash collection in september = (236000*70%*80%) = 132160

So answer is c) $132160

3) January Cash collection = 310000+(934000*20%+934000*80%*75%) = 1057200

So answer is b) $1057200